40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 19 23:27:27 2018, gautam, Update, SUS, ETMX coil driver work in AM tomorrow 
    Reply  Thu Jun 21 00:14:01 2018, gautam, Update, SUS, ETMX coil driver out 
       Reply  Thu Jun 21 03:13:37 2018, gautam, Update, SUS, ETMX coil driver noise ETMXsticthced.pdfETMXFullSpan.pdfETMXCoilDriver.fil.zip
          Reply  Sun Jun 24 20:02:07 2018, gautam, Update, SUS, Some notes about coil driver noise CoilDriverSchematic.pdfD010001_2k_fastOnly_2.25k.pdfD010001_4k_fastOnly_4.5k.pdf
             Reply  Tue Jun 26 16:28:00 2018, gautam, Update, SUS, Coil driver protoboard characterization IMG_7060.JPGETMXstitchced.pdfETMXfullSpan.pdfPSRR.pdf
       Reply  Thu Jun 21 18:25:57 2018, gautam, Update, SUS, ETMX coil driver re-installed 
Message ID: 13993     Entry time: Thu Jun 21 03:13:37 2018     In reply to: 13992     Reply to this: 14012
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMX coil driver noise 

I decided to take a quick look at the data. Changes made to the ETMX coil driver board:

  1. Fast path series resistances: 400 ohm ---> 2.25 kOhm (= 2x 4.5 kohm in parallel). Measured (with DMM) resistance in all 5 paths varied by less than 3 ohms (~0.2%).
  2. All thick film resistors in signal (fast and bias) paths changed to thin film.
  3. AD797 ---> Op27 for monitor output.
  4. Above-mentioned mon output (30Hz HPF-ed) routed to FP LEMO mon via 100ohm for diagnostic purposes.
  5. 4x Trim-pots in analog path removed. 

I also took the chance to check the integrity of the LM6321 ICs. In the past, a large DC offset on the output pin of these has been indicative of a faulty IC. But I checked all the ICs with a DMM, and saw no anomalies.

Measurement condition was that (i) the Fast input was terminated to ground via 50ohm, (ii) the Bias input was shorted to ground. SR785 was used with G=100 Busby preamp (in which Steve installed new batteries today, for someone had left it on for who knows how long) for making the measurement. The voltage measurement was made at the D-Sub connector on the front panel which would be connected to the Sat. Box, with the coil driver not connected to anything downstream.

Summary of results:

[Attachment #1] - Noise measurement out to 800 Hz. The noise only seems to agree with the LISO model above 300 Hz. Not sure if the low-frequency excess is real or a measurement artefact. Tomorrow, I plan to make an LPF pomona box to filter out the HF pickup and see if the low-frequency characteristics change at all. Need to think about what this corner freq. needs to be. In any case, such a device is probably required to do measurements inside the VEA.

[Attachment #2] - Noise measurement for full SR785 span. The 19.5 kHz harmonics are visible. I have a theory about the origin of these, need to do a couple of more tests to confirm and will make a separate log.

[Attachment #3] - zip of LISO file used for modeling coil driver. I don't have the ASCII art in this, so need to double check to make sure I haven't connected some wrong nodes, but I think it's correct.

Measurements seem to be consistent with LISO model predictions.

*Note: Curves labelled "LISO model ..." are really quad sum of liso pred + busby box noise.

My main finding tonight is: With the increased series resistance (400 ohm ---> 2.25 kohm), LISO modeling tells me that even though the series resistance (Johnson noise) used to dominate the voltage noise at the output to the OSEM, the voltage noise of the LT1125 in the bias path now dominates. Since we are planning to re-design the entire bias path anyways, I am not too worried about this for the moment.

I will upload more details + photos + data + schematic + LISO model breakdown tomorrow to a DCC page


gautam noon 21 June 2018: I was looking at the wrong LISO breakdown curves. So the input stage Op27 voltage noise used to dominate. Now the Bias path LT1125 voltage noise dominates. None of the conclusions are affected... I've uploaded the corrected plots and LISO file here now. 

Attachment 1: ETMXsticthced.pdf  206 kB  Uploaded Thu Jun 21 13:17:25 2018  | Hide | Hide all
ETMXsticthced.pdf
Attachment 2: ETMXFullSpan.pdf  274 kB  Uploaded Thu Jun 21 13:17:35 2018  | Hide | Hide all
ETMXFullSpan.pdf
Attachment 3: ETMXCoilDriver.fil.zip  986 Bytes  Uploaded Thu Jun 21 13:18:29 2018
ELOG V3.1.3-